OpenStudy (anonymous):

Suppose Q(1, -1/2) is the midpoint of the line segment AB and suppose A is the point (6,5). Find b.

OpenStudy (anonymous):

Try plotting the points. It'll be easier to visualize. 6-1=5 ... so the x of point B would be 1-5=-4 5-(-1/2)=5.5...so the y of point B would be -1/2-5.5=-6 B (-4,-6)

OpenStudy (anonymous):

Thank you! I've been trying to figure it out for a while now lol